# PrunedLoRA: Robust Gradient-Based structured pruning for Low-rank Adaptation in Fine-tuning

### v1

@@ Abstract
Low-rank adaptation (LoRA) has become a widely used paradigm for parameter-efficient fine-tuning of large language models, yet its representational capacity often lags behind full fine-tuning. Within the context of LoRA, a key open question is how to obtain expressive low-rank adapters from over-parameterized spaces. We propose PrunedLoRA, a new framework that leverages structured pruning to obtain highly representative low-rank adapters from an over-parameterized initialization. Unlike prior approaches that impose a fixed low-rank budget, PrunedLoRA dynamically prunes less important components during fine-tuning and prevents their reactivation, enabling flexible and adaptive rank allocation. For structured pruning, by minimizing the pruning error for overall loss, we provide fine-grained pruning and recovery updates in a gradient-based pruning strategy with grounded interpretation. We provide the first theoretical analysis of the robustness of structured pruning and provably show that under the impact of weight perturbation, gradient-based pruning is more robust than activation-based pruning with respect to overall loss. Empirically, PrunedLoRA consistently outperforms LoRA and its variants across supervised fine-tuning tasks in mathematical reasoning, code generation, and natural language understanding, and it also demonstrates advantages over existing structured pruning methods across diverse sparsity levels.
@@ 1 Introduction
  • We conduct extensive experiments across supervised fine-tuning tasks spanning mathematical reasoning, code generation, and natural language understanding, showing that PrunedLoRA can further narrow the gap between LoRA and FFT. Across different sparsity levels from $50\%$ to $93\%$ and across various pruning tasks (including both dynamic and one-shot pruning), our method consistently outperforms existing structured pruning methods.
@@ 2 Related Work
Low-rank adaptation (LoRA) has been extensively investigated in foundation models [5, 71, 2], with numerous variants and enhancements proposed [48, 23, 69]. [26] assumes that the fine-tuning update can be effectively captured in a low-rank subspace. Specifically, for a pre-trained model with weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, LoRA reparameterizes the weight update $\Delta{\bm{W}}$ via a low-rank decomposition as ${\bm{W}}_{0}+\Delta{\bm{W}}={\bm{W}}_{0}+s{\bm{B}}{\bm{A}}$, where ${\bm{B}}\in\mathbb{R}^{m\times r}$, ${\bm{A}}\in\mathbb{R}^{r\times n}$ and $s=\frac{\alpha}{r}$ is a scaling factor. Here, $r\ll\textit{min}(m,n)$ is the rank of the update. AdaLoRA [89] dynamically allocates the parameter budget by assigning more capacity to task-critical modules, but remains constrained within a limited subspace and does not fully explore the parameter space as in full fine-tuning. LoRA-Prune [88] leverages gradients from LoRA modules rather than the entire model to prune the whole model, which differs from our goal and leads to substantial performance degradation. In contrast, we only prune the trainable parameters to produce representative low-rank adaptations at the end.
@@ 3.1 Motivation
Figure 1: Performance of standard LoRA [26] on GSM8K [7] and HumanEval [4] with different ranks compared to full fine-tuning. Note that the method of full fine-tuning does not involve the initial rank, and we draw a red line here solely for comparison.
Motivation 1: Higher rank results in better performance. As illustrated in Figure 1, employing higher ranks in LoRA consistently leads to improved empirical performance on both GSM8K and HumanEval (see Sec. 4 for details). Notably, as the rank increases, the performance gradually converges toward that of full fine-tuning. This observation motivates our approach: rather than fixing LoRA to a small rank at the outset, we initialize with a sufficiently large rank—providing a number of trainable parameters close to full fine-tuning—and then progressively prune it to a smaller rank. Such a strategy may preserve most of the performance gains in over-parameterized settings while ultimately producing a memory-efficient low-rank adaptation.
@@ 3.2 The Robustness of Gradient-based structured pruning
Activation-based v.s. Gradient-based structured pruning. Pruning induces perturbations to the weights across layers of large language models, which in turn modifies the overall loss and may lead to a deterioration of empirical performance [16, 78]. Within the context of structured pruning [44, 52, 11], activation-based solving Problem (1) and gradient-based pruning using important scores in (2) are two main lines of approaches to find the optimal pruned structure. Intuitively, gradient-based methods focus more on the global correlation [53], so they shall be more robust for the overall loss under the influence of weight perturbation. However, no theoretical analysis provably shows the insight. Here, we analyze the influence of different pruning strategies on the overall loss. We provide formal analysis and general discussion in Appendix B.
Proposition 1 (Unofficial Statement).
@@ 4 Experiment
In this section, we present extensive experiments to evaluate the effectiveness of PrunedLoRA across various tasks and models. With different levels of pruning sparsity, we assess its capabilities on supervised fine-tuning tasks in mathematical reasoning, code generation using the Llama-3-8B model [17], and natural language understanding on a T5-based model covered in Sec 4.1. Then we conduct ablation studies for the hyperparameters, pruning schedules, and more pruning baselines in Sec 4.2 and Appendix C.4. In addition to conducting structured pruning to obtain low-rank adaptation in fine-tuning, one-shot pruning for compressing a pretrained model is crucial in the pre-LLM era [60] as well, but most of the work [19, 60, 16] is activation-based methods without awareness of the influence of weight perturbation on the overall loss function. We provide a simple gradient-based method as well in Appendix D without weight update. It supports the effectiveness of gradient-based pruning strategies for eliminating the impact of weight perturbation.
@@
In addition to fine-tuning baselines, we also compare against existing structured pruning approaches for low-rank adaptation. Gradient-based pruning includes our method, which jointly optimizes parameter updates and pruning structure, as well as the widely used importance-score pruning strategy (Eq. 2) employed in LLM-Pruner [47]. Activation-based pruning determines the pruning structure based on input activation statistics (Eq. 1), as exemplified by ZipLM [32] and SparseGPT [16]. We further include comparisons with other classical pruning strategies [60, 19], along with one-shot pruning, which are reported in Appendix C.4.
@@ 4.1 Experiments on Supervised Fine-tuning
Model and Datasets. To evaluate the scalability of PrunedLoRA, we fine-tune Llama-3-8B on mathematical reasoning and code generation. Besides, we fine-tune a T5-based model on a natural language understanding task. The experimental setup in this study follows closely the protocols established in prior LoRA research [70, 69].
@@
Code: We train our model on a 100k subset of Code-Feedback [94], a high-quality code instruction dataset, removing explanations after code blocks. The model is tested on HumnaEval [4], which consists of 180 Python tasks, and we report the PASS@1 metric.
Beyond the two natural language generation tasks, we further evaluate natural language understanding by fine-tuning a T5-base model [56] on a subset of the GLUE benchmark [68], including MNLI, SST-2, CoLA, QNLI, and MRPC. Model performance is assessed using accuracy as the evaluation metric.
@@
Table 1: Performance comparison of fine-tuning and pruning baselines on GSM8K and HumanEval benchmarks for Llama-3-8B-Base Model. Bold indicates the best result, underline represents the second-best one. ($\uparrow$:  higher values indicate better performance)
@@
Results on Natural Language Generation. Table 1 shows that PrunedLoRA outperforms on both GSM8K and HumanEval. Compared with vanilla LoRA, which lags far behind full fine-tuning (64.4 vs. 73.3 on GSM8K), PrunedLoRA substantially closes the gap, and with init $r=512$ it even matches or surpasses full fine-tuning (73.38 on GSM8K, 48.32 on HumanEval). Relative to structured pruning baselines such as SparseGPT and LLM-Pruner, our method consistently yields higher accuracy, indicating greater robustness. We also find that larger initialization ranks lead to better outcomes, confirming our motivation that starting from higher-rank spaces provides richer expressiveness before pruning down to the final budget. We further provide a more detailed comparison across different initialization ranks for each pruning strategy (SparseGPT, LLM-Pruner, and PrunedLoRA) in Table 5 (See Appendix 5). These results confirm that PrunedLoRA consistently benefits from pruning higher-rank initializations.
@@ 4.2 Experiments on Ablation Study
Pruning Schedule. We also vary the pruning interval ($K1$) and the number of columns pruned per step ($K2$) in Table 6. Gradual pruning with moderate intervals is consistently superior: pruning every 10 steps with $K2$ = 2 achieves the highest accuracy, while aggressive pruning ($K2$ = 4) slightly hurts performance. This suggests that maintaining stability during rank reduction is critical. Besides gradually pruning in post-training, we can also train LoRA with a high rank to converge and do one-shot structure pruning to obtain a low-rank adaptation (Appendix C.4).
Target Rank. Beyond the default rank budget 64 in LoRA, we also examine more aggressive compression (e.g., pruning to target rank in $\{8,16\}$). As expected, extreme pruning leads to performance degradation, but PrunedLoRA remains competitive with or better than activation-based and simple gradient-based baselines at the same target rank (see Appendix C.3). This highlights the robustness of structured pruning with the awareness of the overall under the cases of extreme compression.

### v2

@@ Abstract
Low-rank adaptation (LoRA) has become a widely used paradigm for parameter-efficient fine-tuning of large language models, yet its representational capacity often lags behind full fine-tuning. Within the context of LoRA, a key open question is how to obtain expressive low-rank adapters from over-parameterized spaces. We propose PrunedLoRA, a new framework that leverages structured pruning to obtain highly representative low-rank adapters from an over-parameterized initialization. Unlike prior approaches that impose a fixed low-rank budget, PrunedLoRA dynamically prunes less important components during fine-tuning and prevents their reactivation, enabling flexible and adaptive rank allocation. For structured pruning, by minimizing the pruning error for overall loss, we provide fine-grained pruning and recovery updates in a gradient-based pruning strategy with grounded interpretation. We provide the first theoretical analysis of the robustness of structured pruning and provably show that under the impact of weight perturbation, gradient-based pruning is more robust than activation-based pruning with respect to overall loss. Empirically, PrunedLoRA consistently outperforms LoRA and its variants across supervised fine-tuning tasks in mathematical reasoning and natural language understanding, and it also demonstrates advantages over existing structured pruning methods across diverse sparsity levels.
@@ 1 Introduction
  • We conduct extensive experiments across supervised fine-tuning tasks spanning mathematical reasoning and natural language understanding, showing that PrunedLoRA can further narrow the gap between LoRA and FFT. Across different sparsity levels from $50\%$ to $93\%$ and across various pruning tasks (including both dynamic and one-shot pruning), our method consistently outperforms existing structured pruning methods.
@@ 2 Related Work
Low-rank adaptation (LoRA) has been extensively investigated in foundation models [4, 73, 40, 2, 96], with numerous variants and enhancements proposed [49, 23, 71]. [26] assumes that the fine-tuning update can be effectively captured in a low-rank subspace. Specifically, for a pre-trained model with weight matrix ${\bm{W}}_{0}\in\mathbb{R}^{m\times n}$, LoRA reparameterizes the weight update $\Delta{\bm{W}}$ via a low-rank decomposition as ${\bm{W}}_{0}+\Delta{\bm{W}}={\bm{W}}_{0}+s{\bm{B}}{\bm{A}}$, where ${\bm{B}}\in\mathbb{R}^{m\times r}$, ${\bm{A}}\in\mathbb{R}^{r\times n}$ and $s=\frac{\alpha}{r}$ is a scaling factor. Here, $r\ll\textit{min}(m,n)$ is the rank of the update. AdaLoRA [92] dynamically allocates the parameter budget by assigning more capacity to task-critical modules, but remains constrained within a limited subspace and does not fully explore the parameter space as in full fine-tuning. LoRA-Prune [91] leverages gradients from LoRA modules rather than the entire model to prune the whole model, which differs from our goal and leads to substantial performance degradation. In contrast, we only prune the trainable parameters to produce representative low-rank adaptations at the end.
@@ 3.1 Motivation
Figure 1: Performance of standard LoRA [26] on GSM8K [7] with different ranks compared to full fine-tuning. Note that the method of full fine-tuning does not involve the initial rank, and we draw a red line here solely for comparison.
Motivation 1: Higher rank results in better performance. As illustrated in Figure 1, employing higher ranks in LoRA consistently leads to improved empirical performance on both GSM8K (see Sec. 4 for details). Notably, as the rank increases, the performance gradually converges toward that of full fine-tuning. This observation motivates our approach: rather than fixing LoRA to a small rank at the outset, we initialize with a sufficiently large rank—providing a number of trainable parameters close to full fine-tuning—and then progressively prune it to a smaller rank. Such a strategy may preserve most of the performance gains in over-parameterized settings while ultimately producing a memory-efficient low-rank adaptation.
@@ 3.2 The Robustness of Gradient-based structured pruning
Activation-based v.s. Gradient-based structured pruning. Pruning induces perturbations to the weights across layers of large language models, which in turn modifies the overall loss and may lead to a deterioration of empirical performance [16, 80]. Within the context of structured pruning [45, 53, 11], activation-based solving Problem (1) and gradient-based pruning using important scores in (2) are two main lines of approaches to find the optimal pruned structure. Intuitively, gradient-based methods focus more on the global correlation [54], so they shall be more robust for the overall loss under the influence of weight perturbation. However, no theoretical analysis provably shows the insight. Here, we analyze the influence of different pruning strategies on the overall loss. We provide formal analysis and general discussion in Appendix 7.
Proposition 1 (Unofficial Statement)
@@ 4 Experiment
In this section, we present extensive experiments to evaluate the effectiveness of PrunedLoRA across various tasks and models. With different levels of pruning sparsity, we assess its capabilities on supervised fine-tuning tasks in mathematical reasoning using the Llama-3-8B model [17], and natural language understanding on a T5-based model covered in Sec 4.1. Then we conduct ablation studies for the hyperparameters, pruning schedules, and more pruning baselines in Sec 4.2 and Appendix 8.4. In addition to conducting structured pruning to obtain low-rank adaptation in fine-tuning, one-shot pruning for compressing a pretrained model is crucial in the pre-LLM era [61] as well, but most of the work [19, 61, 16] is activation-based methods without awareness of the influence of weight perturbation on the overall loss function. We provide a simple gradient-based method as well in Appendix 9 without weight update. It supports the effectiveness of gradient-based pruning strategies for eliminating the impact of weight perturbation.
@@
In addition to fine-tuning baselines, we also compare against existing structured pruning approaches for low-rank adaptation. Gradient-based pruning includes our method, which jointly optimizes parameter updates and pruning structure, as well as the widely used importance-score pruning strategy (Eq. 2) employed in LLM-Pruner [48]. Activation-based pruning determines the pruning structure based on input activation statistics (Eq. 1), as exemplified by ZipLM [32] and SparseGPT [16]. We further include comparisons with other classical pruning strategies [61, 19], along with one-shot pruning, which are reported in Appendix 8.4.
@@ 4.1 Experiments on Supervised Fine-tuning
Model and Datasets. To evaluate the scalability of PrunedLoRA, we fine-tune Llama-3-8B on mathematical reasoning. Besides, we fine-tune a T5-based model on a natural language understanding task. The experimental setup in this study follows closely the protocols established in prior LoRA research [72, 71].
@@
Beyond the mathematical reasoning task, we further evaluate natural language understanding by fine-tuning a T5-base model [57] on a subset of the GLUE benchmark [70], including MNLI, SST-2, CoLA, QNLI, and MRPC. Model performance is assessed using accuracy as the evaluation metric.
@@
Table 1: Performance comparison of fine-tuning and pruning baselines on GSM8K benchmark for Llama-3-8B-Base Model. Bold indicates the best result, underline represents the second-best one. ($\uparrow$:  higher values indicate better performance)
@@
Results on Natural Language Generation. Table 1 shows that PrunedLoRA outperforms on both GSM8K. Compared with vanilla LoRA, which lags far behind full fine-tuning (64.4 vs. 73.3 on GSM8K), PrunedLoRA substantially closes the gap, and with init $r=512$ it even matches or surpasses full fine-tuning (73.38 on GSM8K). Relative to structured pruning baselines such as SparseGPT and LLM-Pruner, our method consistently yields higher accuracy, indicating greater robustness. We also find that larger initialization ranks lead to better outcomes, confirming our motivation that starting from higher-rank spaces provides richer expressiveness before pruning down to the final budget. We further provide a more detailed comparison across different initialization ranks for each pruning strategy (SparseGPT, LLM-Pruner, and PrunedLoRA) in Table 5 (See Appendix 5). These results confirm that PrunedLoRA consistently benefits from pruning higher-rank initializations.
@@ 4.2 Experiments on Ablation Study
Pruning Schedule. We also vary the pruning interval ($K1$) and the number of columns pruned per step ($K2$) in Table 6. Gradual pruning with moderate intervals is consistently superior: pruning every 10 steps with $K2$ = 2 achieves the highest accuracy, while aggressive pruning ($K2$ = 4) slightly hurts performance. This suggests that maintaining stability during rank reduction is critical. Besides gradually pruning in post-training, we can also train LoRA with a high rank to converge and do one-shot structure pruning to obtain a low-rank adaptation (Appendix 8.4).
Target Rank. Beyond the default rank budget 64 in LoRA, we also examine more aggressive compression (e.g., pruning to target rank in $\{8,16\}$). As expected, extreme pruning leads to performance degradation, but PrunedLoRA remains competitive with or better than activation-based and simple gradient-based baselines at the same target rank (see Appendix 8.3). This highlights the robustness of structured pruning with the awareness of the overall under the cases of extreme compression.